OpenStudy (anonymous):

A seller allows a discount of 5% on a watch. If he allows a discount of 7%, he earns Rs. 15 less in the profit. What is the marked price of the watch?

OpenStudy (anonymous):

93/100x=15

OpenStudy (anonymous):

93/100x=x-15

ganeshie8 (ganeshie8):

he earns 15rs less if he increases the dicsount to 7%, right ?

ganeshie8 (ganeshie8):

he earns 15rs less if he increases the dicsount from 5% to 7% that means 15rs represents 2% of the cost of watch, right ?

ganeshie8 (ganeshie8):

2% = 15rs 100% = ?

OpenStudy (anonymous):

750